OpenStudy (anonymous):

Whats the square root of 16 and how do you find it?

OpenStudy (uh60blackhawk):

use the calculator to find the square root

OpenStudy (anonymous):

Which button?

OpenStudy (anonymous):

Is it 4?

OpenStudy (uh60blackhawk):

yes it is.

OpenStudy (uh60blackhawk):

do you need anything else

OpenStudy (anonymous):

No thank you!

OpenStudy (uh60blackhawk):

well hope you have nice day and welcome to open study.

OpenStudy (anonymous):

the square root of 16 is 4 because 4*4=16 most people just memorize the perfect squares: 2*2=4 3*3=9 4*4=16 5*5=25 6*6=36 7*7=49 8*8=64 9*9=81 10*10=100 usually you'll only be expected to use calculators for things that involve decimal answers or larger numbers